# Momentum–mass normalized dark–bright solitons to one dimensional Gross–Pitaevskii systems

### v2

@@ Abstract
We rigorously establish the existence of dark-bright solitons as traveling wave solutions to a one dimensional defocusing Gross–Pitaevskii system, a widely used model for describing mixtures of Bose–Einstein condensates and nonlinear optical systems. These solitons are shown to exhibit symmetry and radial monotonicity in modulus, and to propagate at subsonic speed. Our method relies on minimizing an energy functional subject to two constraints: the mass of the bright component and a modified momentum of the dark component. The compactness of minimizing sequences is obtained via a concentration-compactness argument, which requires some novel estimates based on symmetric decreasing rearrangements.
Keywords: Defocusing Schrödinger system, Gross–Pitaevskii system, dark-bright solitons, non standard conditions at infinity, concentration-compactness, symmetric decreasing rearrangements.
@@ 1 Introduction
where $\Psi$ and $\Phi$ are complex-valued functions, and $\alpha,\beta$ are real parameters. This system consists of two coupled nonlinear Schrödinger equations in one spatial dimension and is commonly referred to as the Gross-Pitaevskii system. This is a classical model for describing the dynamics of quasi-one dimensional Bose-Einstein condensates (BECs) consisting of a mixture of two different species [20, 3], as well as for the multi-mode propagation of pulses in nonlinear optical fibers [23, 22]. The sign of $\alpha$ determines the nature of the mutual interactions between species: for $\alpha>0$, the interaction is repulsive (defocusing), while for $\alpha<0$, it is attractive (focusing). Similarly, $\beta>0$ corresponds to self-defocusing interactions in the second component, and $\beta<0$ to self-focusing ones. The case $\beta=0$ is special, as the second equation becomes linear in $\Phi$. In this regime, the system is sometimes referred to as the Gross–Clark system, which models the motion of an uncharged impurity (represented by the wave function $\Phi$) within a Bose-Einstein condensate [18].
In both BECs and optics, coherent structures know as solitons have been experimentally observed. These are localized waves that maintain their shape while traveling at a constant velocity. Based on their qualitative properties, solitons are formally classified into different types. For instance, localized density bumps that vanish at infinity are called bright solitons, while localized density dips with nonzero background are referred to as dark solitons.
Solitonic pairs consisting of combinations of dark and bright components (and other structures such as antidark solitons or domain walls) have also been observed in mixtures of BECs and nonlinear optical systems, though their specific configurations depend on the nature of the internal and mutual interactions [21]. For instance, it is well known that bright solitons do not exist in the scalar $(\alpha=0)$ defocusing $(\beta>0)$ case. However, if both internal and mutual interactions are repulsive, i.e. $\alpha>0$ and $\beta>0$, then the existence of dark-bright soliton pairs is physically justified by a symbiotic relationship between the components, that is, the dark component effectively supports the bright one.
In this work, we focus on the mathematical analysis of dark-bright soliton pairs, i.e. solutions to (1.1) that satisfy the conditions at infinity:
@@
for some $c,\lambda\in\mathbb{R}$, where $u$ and $v$ are complex-valued functions which represent traveling waves with velocity $c$. The equations for $u$ and $v$ are:
@@
The system (1.4) with (1.5) is invariant under translations in $x$ and the phase transformations $(u,v)\mapsto(e^{ik}u,e^{i\ell}v)$ for any $k,\ell\in\mathbb{R}$. Moreover, by complex conjugation, we may assume without loss of generality that $c\geq 0$. Also, since the second equation in (1.4) has only real coefficients, we will restrict our attention to solutions whose second component $v$ is real-valued.
@@
corresponds to the traveling wave profile equation for the defocusing Gross-Pitaevskii equation. It is well-known that, for every subsonic speed, i.e. for every $c\in[0,\sqrt{2})$, there exists a unique (up to translations and phase shift) finite-energy dark soliton solution to (1.7), which is explicit, and its stability properties have been extensively studied [25, 13, 5, 6, 19]. We provide more details on this subject in next section. For a general overview, see the survey [4].
@@
Significant efforts, coming mainly from the physics community, have been devoted to obtaining dark-bright solitons as solutions to the system (1.1) in the coupled case $\alpha\not=0$. However, much of the existing work relies on numerical methods or formal expansions [23, 20]. Despite its obvious relevance for theoretical and experimental physics, a rigorous derivation of fully nontrivial dark-bright solitons as solutions to the Gross–Pitaevskii system is missing in the literature. To the best of our knowledge, only partial results are available. For instance, the case $\beta=0$ has been dealt with in [30, 2]. On the other hand, it is well known that the system (1.4) with the choice $\alpha=\beta=1$, referred to as Manakov system [28], is integrable, and explicit fully nontrivial solutions exist [8, 32]. However, the physically relevant range of parameters extends beyond the integrable case. A broader, though still limited, range for $\alpha$ and $\beta$ is considered in the recent work [29], where semi-explicit dark-bright solitons are constructed via a Lagrangian variational method.
@@
The main contribution of this work is an analytical proof of the existence of dark-bright soliton solutions to (1.1) in the whole defocusing regime, i.e. for any value of the parameters $\alpha>0$ and $\beta\geq 0$. Although the solutions we obtain are not explicit, we show that they are symmetric and radially nonincreasing in modulus. Moreover, we will verify that they propagate at subsonic speed. Our approach relies on minimizing an energy functional under two constraints, one for the momentum of the dark component, and another one for the mass of the bright component, in the spirit of [2]. Arguing this way, the constants $c$ and $\lambda$ arise as Lagrange multipliers, one corresponding to each constraint.
@@
The proof of the compactness of minimizing sequences is based on the well-known Concentration-Compactness principle of Lions [26, 27]. This approach requires excluding the potential issues of vanishing and dichotomy. Ruling out dichotomy for functionals of more than one variable is not a trivial task, and the literature in this matter is limited. To address this challenge, the use of symmetric decreasing rearrangements (see Section 3) has been proven effective. Indeed, a refined Pólya–Szegő inequality due to [14] was the key for establishing compactness for system of Klein–Gordon equations. This technique has also been successfully employed in various other contexts, always for solutions to systems whose both components are real-valued and vanish at infinity. These include nonlinear Schrödinger–Korteweg–de Vries systems [1], focusing nonlinear Schrödinger systems [31], and higher-dimensional problems [33, 16, 17].
@@
In the case of dark-bright solitons, the use of symmetric decreasing rearrangements presents a particular challenge, as the dark component does not vanish at infinity. To address this, nontrivial manipulations of the energy and the momentum of the dark component in terms of its lifting are required. The detailed formulation and the statements of the main results are provided in Section 2.
@@
The rest of the paper is organized as follows. First, Section 3 is devoted to the statements of well-known preliminary results on symmetric decreasing rearrangements. After that, Section 4 includes the proof of Theorem 2.3. Next, Section 5 contains the proof of the compactness of minimizing sequences. Finally, we prove Theorem 2.4 in Section 6.
@@
Notation: Unless a complex range is indicated explicitly, the Lebesgue and Sobolev spaces appearing in this work are of real-valued functions. If $w\in L^{r}(\mathbb{R})$ for some $r\in[1,\infty]$, we denote $\|w\|_{r}=\|w\|_{L^{r}(\mathbb{R})}$. Likewise, if $w\in H^{1}(\mathbb{R})$, we write $\|w\|_{1,2}=\|w\|_{H^{1}(\mathbb{R})}$.
@@ 2 Functional setting and main results
for $q\geq 0$, $m\geq 0$. Notice that we consider only real-valued functions for the second component of the pairs in $\mathcal{X}$. We also stress that, if $u\in\mathcal{NE}(\mathbb{R})$ and $|u|<2$, then $|1-|u||\in(0,1)$, which is precisely the interval where $G$ in (2.1) is increasing. This idea will be behind some proofs of Section 4.
@@
Let us first analyze the semitrivial cases $E_{\textup{min}}(0,m)$ and $E_{\textup{min}}(q,0)$. On the one hand, it is simple to see that
@@
and the infimum is never attained if $m>0$. On the other hand, for $m=0$ we have the following well-known result:
@@
$\mathbf{u}_{q}(x)=\sqrt{\frac{2-c^{2}}{c}}\tanh\left(\frac{\sqrt{2-c^{2}}}{2}x\right)-\frac{c}{\sqrt{2}}i.$ (2.6)
@@
Let $\alpha>0$ and $\beta\geq 0$.
@@
    For every $m\geq 0$, the function $q\mapsto E_{\textup{min}}(q,m)$ is nondecreasing in $0,\infty)$. In particular, $E_{\textup{min}}(q,m)\geq 0$ for all $q\in[0,\pi/2)$, $m\geq 0$.
@@
    In particular, for every $q\in[0,\pi/2)$, the function $m\mapsto E_{\textup{min}}(q,m)$ is nonincreasing, and
@@
The properties of $E_{\textup{min}}$ ultimately lead to the compactness of the minimizing sequences and, in turn, to the existence of a finite-energy solution to (1.4)-(1.5). For instance, we highlight that (2.9) plays a central role in guaranteeing the nonvanishing of minimizing sequences. This inequality can be formally interpreted as expressing an energetic advantage of nontrivial bright components over those with vanishing mass.
@@
Let $\alpha>0$, $\beta\geq 0$, $q\in(0,\pi/2)$, and $m>0$. Assume one of the following two conditions:
@@
We also remark that, in the context of mixtures of BECs, the condition $\alpha^{2}<\beta$ corresponds to the so-called miscibility regime [20, 3]. In physical terms, this conditions says that the repulsive forces between the two species in the mixture of condensates are weaker than the repulsive interatomic interactions. In this regime, i.e. under condition (H1), our main result yields the existence of a dark-bright soliton provided that $E_{\textup{min}}(q,m)$ is small enough. As shown by (2.9), a sufficient condition for (H1) is
@@
In the immiscibility regime $\alpha^{2}>\beta$, and in the transition regime $\alpha^{2}=\beta$ (within which the Manakov case falls), our result also follows under condition (H2). Notice that, by virtue of Theorem 2.1, this condition is trivially satisfied for $m=0$. Moreover, as shown by item (i) in Theorem 2.3, $E_{\textup{min}}(q,m)\geq 0$ for all $q\in(0,\pi/2)$ and $m>0$, which implies that (H2) means a smallness requirement for the mass. It follows also from (2.9) in Theorem 2.3 that a sufficient condition for (H2) is
@@
where, recall, $E_{\textup{min}}(q,0)$ is explicit, see Theorem 2.1. In any case, it remains unclear whether conditions (H1) and (H2) are essential or simply technical.
We conclude this section with a brief discussion on the stability of dark-bright solitons. To the best of the author’s knowledge, a rigorous analytical proof of their orbital stability remains an open problem, even in the Gross–Clark $(\beta=0)$ and the Manakov $(\alpha=\beta=1)$ models. A standard strategy for establishing the orbital stability of constrained minimizers is the approach developed in [10]. However, this method is not directly applicable here, as the quantity $p(u)$, which is fixed in our variational framework, is not conserved along the flow. Nevertheless, the variational structure underlying our existence result suggests, at least on a heuristic level, that orbital stability should hold, and we hope our approach lays the groundwork for a future rigorous proof.
@@ 3 Basics of symmetric decreasing rearrangements
In order to be self-contained, let us state some basic definitions and properties of symmetric decreasing rearrangements. More details can be found in [7, 24].
@@
The following result is commonly known as the Hardy-Littlewood inequality.
@@
We also recall the well-known Pólya-Szegő inequality.
@@
We finally state the version in [1] of the key result due to [14], which essentially says that Pólya-Szegő inequality is strict for functions that decompose in the sum of disjoint bubbles.

### v3

@@ Abstract
We rigorously establish the existence of dark–bright solitons as traveling wave solutions to a one dimensional defocusing Gross–Pitaevskii system, a widely used model for describing mixtures of Bose–Einstein condensates and nonlinear optical systems. These solitons are shown to exhibit symmetry and radial monotonicity in modulus, and to propagate at subsonic speed. Our method relies on minimizing an energy functional subject to two constraints: the mass of the bright component and a modified momentum of the dark component. The compactness of minimizing sequences is obtained via a concentration–compactness argument, which requires some novel estimates based on symmetric decreasing rearrangements.
Keywords: Defocusing Schrödinger system, Gross–Pitaevskii system, dark–bright solitons, non standard conditions at infinity, concentration–compactness, symmetric decreasing rearrangements.
@@ 1 Introduction
where $\Psi$ and $\Phi$ are complex–valued functions, and $\alpha,\beta$ are real parameters. This system consists of two coupled nonlinear Schrödinger equations in one spatial dimension and is commonly referred to as the Gross–Pitaevskii system. This is a classical model for describing the dynamics of quasi–one dimensional Bose–Einstein condensates (BECs) consisting of a mixture of two different species [22, 3], as well as for the multi–mode propagation of pulses in nonlinear optical fibers [25, 24]. The sign of $\alpha$ determines the nature of the mutual interactions between species: for $\alpha>0$, the interaction is repulsive (defocusing), while for $\alpha<0$, it is attractive (focusing). Similarly, $\beta>0$ corresponds to self–defocusing interactions in the second component, and $\beta<0$ to self–focusing ones. The case $\beta=0$ is special, as the second equation becomes linear in $\Phi$. In this regime, the system is sometimes referred to as the Gross–Clark system, which models the motion of an uncharged impurity (represented by the wave function $\Phi$) within a Bose–Einstein condensate [20].
In both BECs and optics, coherent structures known as solitons have been experimentally observed. These are localized waves that maintain their shape while traveling at a constant velocity. Based on their qualitative properties, solitons are formally classified into different types. For instance, localized density bumps that vanish at infinity are called bright solitons, while localized density dips with nonzero background are referred to as dark solitons.
Solitonic pairs consisting of combinations of dark and bright components (and other structures such as antidark solitons or domain walls) have also been observed in mixtures of BECs and nonlinear optical systems, though their specific configurations depend on the nature of the internal and mutual interactions [23]. For instance, it is well known that bright solitons do not exist in the scalar $(\alpha=0)$ defocusing $(\beta>0)$ case. However, if both internal and mutual interactions are repulsive, i.e. $\alpha>0$ and $\beta>0$, then the existence of dark–bright soliton pairs is physically justified by a symbiotic relationship between the components, that is, the dark component effectively supports the bright one.
In this work, we focus on the mathematical analysis of dark–bright soliton pairs, i.e. solutions to (1.1) that satisfy the conditions at infinity:
@@
for some $c,\lambda\in\mathbb{R}$, where $u$ and $v$ are complex–valued functions which represent traveling waves with velocity $c$. The equations for $u$ and $v$ are:
@@
The system (1.4) with (1.5) is invariant under translations in $x$ and the phase transformations $(u,v)\mapsto(e^{ik}u,e^{i\ell}v)$ for any $k,\ell\in\mathbb{R}$. Moreover, by complex conjugation, we may assume without loss of generality that $c\geq 0$. Also, since the second equation in (1.4) has only real coefficients, we will restrict our attention to solutions whose second component $v$ is real–valued. In fact, it will be shown in the Appendix that the bright component to system (1.4) is real–valued up to a multiplication by a complex constant of modulus one. Notice that this is no longer the case for other solitonic pairs such as dark–dark solitons, for which both components have nontrivial imaginary part, see [35] for instance.
@@
corresponds to the traveling wave profile equation for the defocusing Gross–Pitaevskii equation. It is well–known that, for every subsonic speed, i.e. for every $c\in[0,\sqrt{2})$, there exists a unique (up to translations and phase shift) finite–energy dark soliton solution to (1.7), which is explicit, and its stability properties have been extensively studied [27, 15, 5, 6, 21]. We provide more details on this subject in next section. For a general overview, see the survey [4].
@@
Significant efforts, coming mainly from the physics community, have been devoted to obtaining dark–bright solitons as solutions to the system (1.1) in the coupled case $\alpha\not=0$. However, much of the existing work relies on numerical methods or formal expansions [25, 22]. Despite its obvious relevance for theoretical and experimental physics, a rigorous derivation of fully nontrivial dark–bright solitons as solutions to the Gross–Pitaevskii system is missing in the literature. To the best of our knowledge, only partial results are available. For instance, the case $\beta=0$ has been dealt with in [32, 2]. On the other hand, it is well known that the system (1.4) with the choice $\alpha=\beta=1$, referred to as Manakov system [30], is integrable, and explicit fully nontrivial solutions exist [8, 35]. However, the physically relevant range of parameters extends beyond the integrable case. A broader, though still limited, range for $\alpha$ and $\beta$ is considered in the recent work [31], where semi–explicit dark–bright solitons are constructed via a Lagrangian variational method.
The main contribution of this work is an analytical proof of the existence of dark–bright soliton solutions to (1.1) in the so–called miscibility regime i.e. for any value of the parameters $\alpha>0$ and $\beta>0$ satisfying $\alpha^{2}\leq\beta$. Moreover, we show that these solutions are symmetric and radially nonincreasing in modulus, and they propagate at subsonic speed. In the context of mixtures of BECs, the condition $\alpha^{2}\leq\beta$ indicates that the repulsive forces between the two species in the mixture of condensates are weaker than the repulsive interatomic interactions [22, 3]. Mathematically, in the miscibility regime, the renormalized energy functional (2.4) defined below is nonnegative, as proved in Theorem 2.3. This property plays a crucial role in establishing the existence result. In contrast, in the immiscibility regime corresponding to $\alpha^{2}>\beta$, the renormalized energy is expected to attain negative values, and the problem therefore requires a different analytical approach. This case will be addressed in future work.
@@
Our approach relies on minimizing the mentioned energy functional under two constraints, one for the momentum of the dark component, and another one for the mass of the bright component, in the spirit of [2]. Arguing this way, the constants $c$ and $\lambda$ arise as Lagrange multipliers, one corresponding to each constraint.
@@
The proof of the compactness of minimizing sequences is based on the well–known Concentration–Compactness principle of Lions [28, 29]. This approach requires excluding the potential issues of vanishing and dichotomy. Ruling out dichotomy for functionals of more than one variable is not a trivial task, and the literature in this matter is limited. To address this challenge, the use of symmetric decreasing rearrangements (see Section 3) has been proven effective. Indeed, a refined Pólya–Szegő inequality due to [16] was the key for establishing compactness for systems of Klein–Gordon equations. This technique has also been successfully employed in various other contexts, always for solutions to systems whose both components are real–valued and vanish at infinity. These include nonlinear Schrödinger–Korteweg–de Vries systems [1], focusing nonlinear Schrödinger systems [34], and higher–dimensional problems [36, 18, 19].
@@
In the case of dark–bright solitons, the use of symmetric decreasing rearrangements presents a particular challenge, as the dark component does not vanish at infinity. To address this, nontrivial manipulations of the energy and the momentum of the dark component in terms of its lifting are required. The detailed formulation and the statements of the main results are provided in Section 2.
@@
The rest of the paper is organized as follows. First, Section 3 is devoted to the statements of well–known preliminary results on symmetric decreasing rearrangements. After that, Section 4 includes the proof of Theorem 2.3. Next, Section 5 contains the proof of the compactness of minimizing sequences. Finally, we prove Theorem 2.4 in Section 6.
@@
Funding: This work has been supported by the Grant PID2024–156079NA–I00 of the MCIU/AEI (Spain).
Notation: Unless a complex range is indicated explicitly, the Lebesgue and Sobolev spaces appearing in this work are of real–valued functions. If $w\in L^{r}(\mathbb{R})$ for some $r\in[1,\infty]$, we denote $\|w\|_{r}=\|w\|_{L^{r}(\mathbb{R})}$. Likewise, if $w\in H^{1}(\mathbb{R})$, we write $\|w\|_{1,2}=\|w\|_{H^{1}(\mathbb{R})}$.
@@ 2 Functional setting and main results
for $q\geq 0$, $m\geq 0$. Notice that we consider only real–valued functions for the second component of the pairs in $\mathcal{X}$. We also stress that, if $u\in\mathcal{NE}(\mathbb{R})$ and $|u|<2$, then $|1-|u||\in(0,1)$, which is precisely the interval where $G$ in (2.1) is increasing. This idea will be behind some proofs of Section 4.
@@
Let us first analyze the semitrivial cases $E_{\textup{min}}(0,m)$ and $E_{\textup{min}}(q,0)$. On the one hand, in the miscible regime $\alpha^{2}\leq\beta$, it is simple to prove (see Theorem 2.3) that
@@
and the infimum is never attained if $m>0$. On the other hand, for $m=0$ we have the following well–known result:
@@
$\mathbf{u}_{q}(x)=\sqrt{\frac{2-c^{2}}{{2}}}\tanh\left(\frac{\sqrt{2-c^{2}}}{2}x\right)-\frac{c}{\sqrt{2}}i.$ (2.6)
@@
Let $\alpha>0$ and $\beta>0$ satisfy $\alpha^{2}\leq\beta$.
@@
    For every $m\geq 0$, the function $q\mapsto E_{\textup{min}}(q,m)$ is nondecreasing in $0,\infty)$ and $E_{\textup{min}}(0,m)=0$. In particular, $E_{\textup{min}}(q,m)\geq 0$ for all $q\in[0,\pi/2)$, $m\geq 0$.
@@
    In particular, for every $q\in[0,\pi/2)$, the function $m\mapsto E_{\textup{min}}(q,m)$ is nonincreasing in $0,\infty)$, and
@@
The properties of $E_{\textup{min}}$ ultimately lead to the compactness of the minimizing sequences and, in turn, to the existence of a finite–energy solution to (1.4)–(1.5). For instance, we highlight that (2.9) plays a central role in guaranteeing the nonvanishing of minimizing sequences. This inequality can be formally interpreted as expressing an energetic advantage of nontrivial bright components over those with vanishing mass.
@@
Let $\alpha>0$ and $\beta>0$ satisfy $\alpha^{2}\leq\beta$, and let $q\in(0,\pi/2)$ and $m>0$. Assume one of the following two conditions:
@@
Some comments about conditions (H1) and (H2) are in order. First, under condition (H1), our main result yields the existence of a dark–bright soliton provided that $E_{\textup{min}}(q,m)$ is small enough. As shown by (2.9), a sufficient condition for (H1) is
@@
Our result also follows under condition (H2). Notice that, by virtue of Theorem 2.1, this condition is trivially satisfied for $m=0$. Moreover, as shown by item (i) in Theorem 2.3, $E_{\textup{min}}(q,m)\geq 0$ for all $q\in(0,\pi/2)$ and $m>0$, which implies that (H2) means a smallness requirement for the mass. It follows also from (2.9) in Theorem 2.3 that a sufficient condition for (H2) is
@@
where, recall, $E_{\textup{min}}(q,0)$ is explicit, see Theorem 2.1.
Regarding the necessity of conditions (H1) and (H2), we remark that they are certainly not necessary for the mere existence of dark–bright solitons to (1.1) in the regime $\alpha^{2}\leq\beta$. Indeed, the Manakov case ($\alpha=\beta=1$) is not covered by our assumptions when $m$ is large, even though explicit dark–bright solitons exist for all $m$, see [8, 35]. However, the Manakov system possesses special features, namely its integrability and its transitioning nature between the miscible and immiscible regimes, so that a general existence proof allowing $\alpha^{2}<\beta$ may in principle require additional hypotheses. In any case, it remains unclear whether conditions (H1) and (H2) are essential for our approach or simply technical.
We conclude this section with a brief discussion on the stability of dark–bright solitons. To the best of the author’s knowledge, a rigorous analytical proof of their orbital stability remains an open problem, even in the Gross–Clark $(\beta=0)$ and the Manakov $(\alpha=\beta=1)$ models. A standard strategy for establishing the orbital stability of constrained minimizers is the approach developed in [10]. However, this method is not directly applicable here, as the quantity $p(u)$, which is fixed in our variational framework, is not conserved along the flow. It is worth noting that the global renormalized momentum
$Q(\Psi)+P(\Phi):=-\frac{1}{2}\int_{\mathbb{R}}\langle i\partial_{x}\Psi,\Psi\rangle\frac{1-|\Psi|^{2}}{|\Psi|^{2}}\,\mathrm{d}x+\frac{1}{2}\int_{\mathbb{R}}\langle i\partial_{x}\Phi,\Phi\rangle\,\mathrm{d}x,$ (2.12)
is indeed a conserved quantity, at least formally. Consequently, in order to prove orbital stability, it would be desirable to establish a variational minimization principle at fixed global renormalized momentum that yields fully nontrivial minimizers. However, our approach based on symmetric decreasing rearrangements does not appear to be well suited to the functional (2.12) as a constraint. Nevertheless, the variational structure underlying our existence result suggests, at least on a heuristic level, that orbital stability should hold, and we hope our approach lays the groundwork for a future rigorous proof.
@@ 3 Basics of symmetric decreasing rearrangements
In order to be self–contained, let us state some basic definitions and properties of symmetric decreasing rearrangements. More details can be found in [7, 26].
@@
The following result is commonly known as the Hardy–Littlewood inequality.
@@
We also recall the well–known Pólya–Szegő inequality.
@@
We finally state the version in [1] of the key result due to [16], which essentially says that Pólya–Szegő inequality is strict for functions that decompose in the sum of disjoint bubbles.